Find all solutions by factoring.
(6x + 60)(x - 17) = (5x + 58)(x - 18)

Answer by mananth(16946)

(6x + 60)(x - 17) = (5x + 58)(x - 18)
6x^2-102x+60x-1020 = 5x^2-90x+58x-1044
x^2-102x+60x+90x-58x-1020+1044=0
x^2-10x+24=0
x^2-6x-4x+24=0
x(x-6)-4(x-6)=0
(x-4)(x-6)=0x=4 Or 6